OpenStudy (anonymous):

ANSWER THIS PLEASE Find the sum of 208+210+212...+422 using Gauss's method

OpenStudy (wolf1728):

You want to sum 208 + 210 + 212 ... 418 + 420 + 422 Gauss (as a kid) summed 1 through 100 by adding the numbers 100 through 1 to the numbers 1 through 100. He realized it would create 100 occurrences of the number 101 divided by 2. So 101 * 100 / 2 = 10,100 / 2 = 5,050 So let's do the same with that problem. Adding 422 420 418 to the series produces 108 occurrences of 630. The sum of the series would be 108*630 / 2 or 34,020

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."